the length of a rectangle is twice the width, its area is 20 square inches. find the dimensions of the rectangle
Answer by TimothyLamb(4379)

L = 2w
Lw = 20
---
Lw = 20
(2w)w = 20
ww = 20/2
w = sqrt( 20/2 )
w = 3.16227766017 inches
L = 6.32455532034 inches
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php